​A stationary man hears a sound of frequency 447 Hz coming from a source moving towards him


A stationary man hears a sound of frequency 447 Hz coming from a source moving towards him. The frequency of sound from the source is 385 Hz. Calculate the speed of the source in units of m/s. Assume v = 340 m/s.
Answer #1

Using doppler's effect:

frequency received by men will be, when Source is moving towards statinary observer:

f1 = f0*(V - Vo)/(V - Vs)

f0 = frequenct of sound = 385 Hz

V = Speed of Sound = 340 m/sec

Vo = speed of Observer(men) = 0 m/sec

Vs = speed of soure = ??

f1 = frequency observed by man = 447 Hz

then, 447 = 385*(340 - 0)/(340 - Vs)

340 - Vs = 385*340/447

Vs = 340 - 385*340/447

Vs = 47.16 m/s

In 3 significant figure:

Vs = speed of the source = 47.2 m/s

"Let me know if you have any query."